# Rui Lia,  Anyao Wangb   Mingqing Zhaic

### v2

@@ Abstract
A graph $G$ is said to satisfy the Hereditarily Bounded Property $P_{t,r}$ if, for every subgraph $H$ of $G$ with at least $t$ vertices, the size of $H$ satisfies $|E(H)|\leq t\cdot|V(H)|+r$. In 1996, Guiduli posed the following problem: If an $n$-vertex graph $G$ possesses the property $P_{t,r}$, where $t$ is a positive real number and $r$ is a real number such that $r\geq-\binom{\lfloor t+1\rfloor}{2}$, does it follow that the spectral radius $\rho(G)$ is bounded above by $c(s,t)+\sqrt{\lfloor t\rfloor n}$, where $s=\binom{t+1}{2}+r$?
In this paper, we affirmatively resolve Guiduli’s problem. Furthermore, we provide a precise characterization of the extremal graphs that achieve this bound. Specifically, these extremal graphs are constructed as the join of two graphs: one being a complete graph $K_{\lfloor t\rfloor}$, and the other either a disjoint union of $K_{3}$ and isolated vertices, or a forest consisting solely of star structures.
The forest of order $n-\lfloor t\rfloor$ in extremal graphs is meticulously characterized. The key to our proof is the classification of the stars into four categories–maximal, large, small, and minimal–according to their sizes. This classification provides deeper insights into how edge distribution influence the spectral properties of graphs under the constraint of property $P_{t,r}$.
Keywords: Hereditarily Bounded Property; spectral radius; extremal graph
@@ 1 Introduction
All graphs considered in this context are simple, undirected and free from isolated vertices. We use $|G|$ and $e(G)$ to denote the numbers of vertices and edges in a graph $G$, respectively. For a subset $S$ of $V(G)$, let $G[S]$ be the subgraph induced by $S$. Let $N_{U}(v)$ denote the set of neighbors of vertex $v$ within the subset $U$ of vertices, formally defined as $N_{U}(v)=\{u\mid u\in U\text{ and }uv\in E(G)\}$. In particular, when $U=V(G)$, we simply write $N(v)$ for $N_{V(G)}(v)$. Denoted by $d_{U}(v)$ the cardinality of the set $N_{U}(v)$. Moreover, let $e(U)$ denote the size of $G[U]$ and $e(U,W)$ be the number of edges connecting $U$ and $W$ in $G$. Let $kG$ be the disjoint union of $k$ copies of $G$, and $G_{1}\nabla G_{2}$ be the graph obtained from $G_{1}$ and $G_{2}$ by joining each vertex of $G_{1}$ with each vertex of $G_{2}$.
@@
In Section 4, we provide a detailed characterization of the forest structure, classifying the stars into maximal, big, small, and minimal stars based on their sizes.
@@ 2 Stability theorem
Let $\mathcal{G}$ denote the family of $n$-vertex graphs that satisfy property $P_{t,r}$. For any $G\in\mathcal{G}$, we assume for simplicity that the graph $G$ considered in this paper is connected. If $G$ were not connected, we could identify a component $C$ within $G$ such that the induced subgraph $G[C]$ also satisfies property $P_{t,r}$ and has the same spectral radius as $G$, i.e., $\rho(G[C])=\rho(G)$. Therefore, it suffices to consider only the connected component $G[C]$ for our analysis. Furthermore, we may assume that $n$ is sufficiently large. If $n$ were a finite constant, we could set $c(s,t)=n$. In this case, the spectral radius would be bounded by $\rho(G)\leq\rho(K_{n})=n-1\leq c(s,t)+\sqrt{\lfloor t\rfloor n}$. Consequently, Theorem 1.2 would hold trivially.
Thus, for the remainder of our discussion, we focus on connected graphs $G\in\mathcal{G}$ with a sufficiently large order $n$. Given that $G\in\mathcal{G}$, it implies that $t\geq 1$. To see why, suppose for contradiction that $10<t<$ 1. For sufficiently large $n$, $e(G)\leq tn+r<n-1$. However, since $G$ is connected, it must have at least $n-1$ edges. This leads to a contradiction, hence $t\geq 1$.
@@
For convenience, we define $s=\big{\lceil}\frac{t(\lfloor t\rfloor+1)+r}{\lfloor t\rfloor+1-t}\big{\rceil}+1$ in this paper. This definition of $s$ differs from that given in Guiduli’s thesis but similarly depends on $t$ and $r$. It is clear that $s>t+1$ as $r\geq-\binom{\lfloor t+1\rfloor}{2}$. To prove our main result, we introduce a stability theorem that plays a crucial role in our analysis.
@@
Let $G$ be a graph of sufficiently large order $n$. Let $X$ be a non-negative eigenvector corresponding to $\rho(G)$, and $x_{u^{*}}=\max_{u\in V(G)}x_{u}$. If $G$ has property $P_{t,r}$ and $\rho(G)\geq\sqrt{\lfloor t\rfloor(n-\lfloor t\rfloor)}$, then there exists a vertex subset $V$ of $G$ with $|V|=\lfloor t\rfloor$ such that $x_{v}\geq(1-\frac{1}{50s^{2}})x_{u^{*}}$ and $d_{G}(v)\geq(1-\frac{1}{25s^{2}})n$ for every $v\in V$.
@@
Given that $\rho(G)\geq\sqrt{\lfloor t\rfloor(n-\lfloor t\rfloor)}$, we have the following inequality for any vertex $u\in V(G)$:
@@
Note that $N(u)=A_{1}\cup B_{1}$. Then $N(v)\setminus\{u\}=N_{A_{1}\cup A_{2}}(v)\cup N_{B_{1}\cup B_{2}}(v)$ for each $v\in N(u)$. By the definition of $V_{1}$, it follows that $x_{w}<(5s)^{-4}x_{u^{*}}$ for every $w\in B_{1}\cup B_{2}$.
Therefore, we can establish an upper bound for the vertices in $A_{1}$ as follows:
@@
Proof of Theorem 2.1: We start by proving lower bounds on $x_{u}$ and $|N(u)|$ for each $u\in V$. Suppose, for the sake of contradiction, that there exists $u_{0}\in V$ such that $x_{u_{0}}<(1-\frac{1}{50s^{2}})x_{u^{*}}.$
@@
To show $|V|=\lfloor t\rfloor$, assume first that $|V|\geq\lfloor t\rfloor+1$. By the previous inequality, every vertex $u\in V$ has at most $\frac{n}{(5s)^{2}}$ non-neighbors. Correspondingly, any $\lfloor t\rfloor+1$ vertices in $V$ have at least $n-(\lfloor t\rfloor+1)\frac{n}{(5s)^{2}}$ common neighbors. Notice that $n-(\lfloor t\rfloor+1)\frac{n}{(5s)^{2}}\geq s$. Then $G$ contains a bipartite subgraph isomorphic to $K_{\lfloor t\rfloor+1,s}$. By a similar argument in the proof of Claim 2.4, there is a contradiction since $K_{\lfloor t\rfloor+1,s}$ does not satisfy the property $P_{t,r}$. Therefore, we conclude $|V|\leq\lfloor t\rfloor.$
@@
Combining both parts, we establish the desired results, $x_{u}\geq\big{(}1-\frac{1}{50s^{2}}\big{)}x_{u^{*}}$ and $|N(u)|\geq\big{(}1-\frac{1}{25s^{2}}\big{)}n$ for each $u\in V$. Moreover, $|V|=\lfloor t\rfloor$. The proof of Theorem 2.1 is now complete. ∎
@@
Choose a vertex $u\in U\cup W$. By Claim 3.1, let $N_{U}(u)=\{u_{1},u_{2},\dots,u_{d}\}$ and denote by $H$ the subgraph in $G$ induced by the set $V\cup N_{U}(u)\cup\{u\}$. Then $e(H)\geq(\lfloor t\rfloor+1)d$. Since $H$ has the property $P_{t,r}$, we have $e(H)\leq t|H|+r=t(\lfloor t\rfloor+1+d)+r$. Combining the two inequalities, we get that $(\lfloor t\rfloor+1-t)d\leq t(\lfloor t\rfloor+1)+r$ and so
$d_{U}(u)=d\leq\frac{t(\lfloor t\rfloor+1)+r}{\lfloor t\rfloor+1-t}\leq s-1.$
Recall that $W$ is not complete to $V$, meaning each vertex in $W$ is not adjacent to some vertex in $V$. By Theorem 2.1, every vertex in $V$ is not adjacent to at most $\frac{n}{25s^{2}}$ vertices of $G$. Therefore, we have $|W|\leq\frac{\lfloor t\rfloor n}{25s^{2}}$. Note that $d(u)\leq d_{U}(u)+|V|+d_{W}(u)\leq s-1+\lfloor t\rfloor+d_{W}(u)$. Moreover, since $G$ has the property $P_{t,r}$, we have either $e(W)\leq t|W|+r$ or $|W|<t$. Consequently, for $w\in W$, we get that
@@
Now, suppose that $W\neq\emptyset$. Let $G^{*}$ be a graph obtained from $G$ by
  • Removing all edges within $G[W]$.
  • Removing all edges between $U$ and $W$.
  • Adding edges between $V\setminus N(W)$ and $W$.
Additionally, let $w_{1}$ be a vertex of minimal degree in $G[W]$, and let $E_{1}$ be the set of edges incident to $w_{1}$ in $G[W]$. Following the same approach, let $w_{2}$ be a vertex of minimal degree in $G[W\setminus\{w_{1}\}]$, and let $E_{2}$ be the set of edges incident to $w_{2}$ in $G[W\setminus\{w_{1}\}]$. Repeating the process, we define $W_{1}=W$ and $W_{i}=W\setminus\{w_{1},\dots,w_{i-1}\}$ for $i\geq 2$, where $w_{i}$ is a vertex of minimum degree in $G[W_{i}]$ and $E_{i}$ is the set of edges that are incident to $w_{i}$ in $G[W_{i}]$. By continuing this process iteratively, we have
@@
In view of the definition of $W$, we have $|V\setminus N_{V}(w_{i})|\geq 1$ for each $w_{i}\in W$. By Theorem 2.1, $x_{v}\geq\big{(}1-\frac{1}{50s^{2}}\big{)}x_{u^{*}}$ for each $v\in V$. Moreover, by previous discussion, we know that $d_{U}(u)\leq s-1$ for each $u\in U$ and $d_{W_{i}}(w_{i})<2s$ for each $w_{i}\in W_{i}$. Combining (10) and (3), we can see that
@@
By the maximality of $G$, the modified graph $G^{*}$ cannot satisfy the property $P_{t,r}$. Specifically, there exists an induced subgraph $H$ of $G^{*}$ such that $|H|\geq t$ and $e(H)>t|H|+r$. Since $G^{*}[U\cup V]=G[U\cup V]$, the violation of the property $P_{t,r}$ must involve vertices in $W$, i.e., $V(H)\cap W\neq\emptyset$.
Let $H^{\prime}=G^{*}[V(H)\setminus W]$. According to the definition of $G^{*}$, the size of $H^{\prime}$ can be expressed as: $e(H)=e(H^{\prime})+|V(H)\cap W|\cdot|V(H)\cap V|>t|V(H)|+r.$ Since $|V(H)\cap V|\leq\lfloor t\rfloor\leq t$, it follows that $e(H^{\prime})>t|V(H)|+r-t|V(H)\cap W|=t|V(H^{\prime})|+r$.
If $|V(H^{\prime})|\geq t$, this inequality leads to a contradiction because $G^{*}[V(H^{\prime})]=G[V(H^{\prime})]$, and $G$ is assumed to satisfy the property $P_{t,r}$.
If $|V(H^{\prime})|<t$, given that $|U\setminus V(H^{\prime})|\geq n/2-t\geq|W|$, there are sufficiently many vertices in $U$ to replace $V(H)\cap W$. We can define $H^{\prime\prime}=G^{*}[U^{\prime}\cup V(H^{\prime})]$, where $U^{\prime}\subseteq U\setminus V(H^{\prime})$ and $|U^{\prime}|=|W|$. Now, the size of $H^{\prime\prime}$ can be expressed as: $e(H^{\prime\prime})=e(H)+e(V(H)\cap U,U^{\prime})+e(U^{\prime})>t|V(H)|+r=t|V(H^{\prime})|+r$. This also leads to a contradiction.
@@
By Claim 3.2, we have $V(G)=V\cup U$ and $e(V,U)=|V||U|$, where $|V|=\lfloor t\rfloor$. Recall that every $u\in U$ has at most $s-1$ neighbors in $U$. Let $u^{\star}\in U$ such that $x_{u^{\star}}=\max_{u\in U}x_{u}$. Then we have $\rho x_{u^{\star}}\leq(s-1)x_{u^{\star}}+\lfloor t\rfloor x_{u^{*}}.$ It follows that $x_{u^{\star}}\leq\frac{\lfloor t\rfloor}{\rho-s+1}x_{u^{*}}.$ Also, since $u^{*}\in V$, we have $\rho x_{u^{*}}\leq(\lfloor t\rfloor-1)x_{u^{*}}+(n-\lfloor t\rfloor)x_{u^{\star}}.$ Combining these two inequalities, we get that
@@
4 Extremal graphs
In this section, we will prove Theorem 1.3 and characterize the structure of the extremal graph. Before delving into our proof, we first introduce some concepts and some key lemmas. For a positive integer $\ell$, an $\ell$-walk in a graph $G$ consists of $\ell+1$ vertices where consecutive vertices are adjacent. Let $w_{G}^{\ell}(u)$ be the number of $\ell$-walks starting at vertex $u$ in $G$. The total number of $\ell$-walks in $G$ is defined as $W_{\ell}(G)=\sum_{v\in V(G)}w_{G}^{\ell}(v)$. Obviously, $W_{1}(G)=2e(G)$ and $W_{2}(G)=\sum_{v\in V(G)}(d_{G}(v))^{2}.$
Let $G_{1}$ and $G_{2}$ be two graphs, possibly of different orders. We say $G_{1}$ is strictly walk-preferable to $G_{2}$, if there exists a positive integer $\ell$ such that $W_{\ell}(G_{1})>W_{\ell}(G_{2})$ and $W_{i}(G_{1})=W_{i}(G_{2})$ for every $i\leq\ell-1$. Otherwise, $W_{\ell}(G_{1})=W_{\ell}(G_{2})$ for all $\ell\geq 1$, then we say $G_{1}$ is walk-equivalent to $G_{2}$ and denote $G_{1}\equiv G_{2}$. We write $G_{2}\prec G_{1}$ if $G_{1}$ is strictly walk-preferable to $G_{2}$, and we write $G_{2}\preceq G_{1}$ if $G_{1}$ is walk-preferable to $G_{2}$ (either strictly or equivalently).

### v3

# Rui Lia,  Anyao Wangb,   Mingqing Zhaic

@@ Abstract
We resolve a problem posed by Guiduli (1996) on the spectral radius of graphs satisfying the Hereditarily Bounded Property $P_{t,r}$, which requires that every subgraph $H$ with $|V(H)|\geq t$ satisfies $|E(H)|\leq t|V(H)|+r$. For an $n$-vertex graph $G$ satisfying $P_{t,r}$, where $t>0$ and $r\geq-\binom{\lfloor t+1\rfloor}{2}$, we prove that the spectral radius $\rho(G)$ is bounded above by $\rho(G)\leq c(s,t)+\sqrt{\lfloor t\rfloor n}$, where $s=\binom{\lfloor t\rfloor+1}{2}+r$, thus affirmatively answering Guiduli’s conjecture.
Furthermore, we present a complete characterization of the extremal graphs that achieve this bound. These graphs are constructed as the join graph $K_{\lfloor t\rfloor}\nabla F$, where $F$ is either $K_{3}\cup(n-\lfloor t\rfloor-3)K_{1}$ or a forest consisting solely of star structures. The specific structure of such forests is meticulously characterized. Central to our analysis is the introduction of a novel potential function $\eta(F)=e(F)+(\lfloor t\rfloor-t)|V(F)|$, which quantifies the structural ”positivity” of subgraphs. By combining edge-shifting operations with spectral radius maximization principles, we establish sharp bounds on $\eta^{+}(G)$, the cumulative positivity of $G$. Our results contribute to the understanding of spectral extremal problems under edge-density constraints and provide a framework for analyzing similar hereditary properties.
Keywords: Hereditarily Bounded Property; spectral radius; extremal graph; edge-shifting
@@ 1 Introduction
All graphs considered in this context are simple, undirected and free from isolated vertices. We use $|G|$ and $e(G)$ to denote the numbers of vertices and edges in a graph $G$, respectively. For a subset $S$ of $V(G)$, let $G[S]$ be the subgraph induced by $S$. Let $N_{U}(v)$ denote the set of neighbors of vertex $v$ within the subset $U$ of vertices, formally defined as $N_{U}(v)=\{u\mid u\in U\text{ and }uv\in E(G)\}$. In particular, when $U=V(G)$, we simply write $N(v)$ for $N_{V(G)}(v)$. Denoted by $d_{U}(v)$ the cardinality of the set $N_{U}(v)$. Moreover, let $e(U)$ denote the size of $G[U]$ and $e(U,W)$ be the number of edges connecting $U$ and $W$ in $G$. Let $kG$ be the disjoint union of $k$ copies of $G$, and $G_{1}\nabla G_{2}$ be the graph obtained by joining each vertex of $G_{1}$ with each vertex of $G_{2}$.
@@
In Section 4, we provide a detailed characterization of the forest structure in Theorem 1.3, classifying the stars into maximal, big, small, and minimal stars based on their sizes. Therefore, all the extremal graphs are completely characterized.
@@ 2 Stability theorem
Let $\mathcal{G}$ denote the family of $n$-vertex graphs that satisfy property $P_{t,r}$. For any $G\in\mathcal{G}$, we assume for simplicity that the graph $G$ considered in this paper is connected. If $G$ were not connected, we could identify a component $C$ within $G$ such that the induced subgraph $G[C]$ also satisfies property $P_{t,r}$ and has the same spectral radius as $G$, i.e., $\rho(G[C])=\rho(G)$. Therefore, it suffices to consider only the connected component $G[C]$ for our analysis. Furthermore, we may assume that $n$ is sufficiently large. If $n$ were a finite constant, we could set $c(s,t)=n$. In this case, the spectral radius would be bounded by $\rho(G)\leq\rho(K_{n})=n-1\leq c(s,t)+\sqrt{\lfloor t\rfloor n}$. Hence, Theorem 1.2 would hold trivially.
Thus, for the remainder of our discussion, we focus on connected graphs $G\in\mathcal{G}$ with a sufficiently large order $n$. Given that $G\in\mathcal{G}$, it implies that $t\geq 1$. To see why, suppose for contradiction that $10<t<$ 1. For sufficiently large $n$, $e(G)\leq tn+r<n-1$. However, since $G$ is connected, $e(G)\geq n-1$. This leads to a contradiction, hence $t\geq 1$.
@@
For convenience, we define $s=\big{\lceil}\frac{t(\lfloor t\rfloor+1)+r}{\lfloor t\rfloor+1-t}\big{\rceil}+1$ in this paper. This definition of $s$ differs from that given in Guiduli’s thesis but similarly depends on $t$ and $r$. It is clear that $s>t+1$ as $r\geq-\binom{\lfloor t\rfloor+1}{2}$. To prove our main result, we introduce a stability theorem that plays a crucial role in our analysis.
@@
Let $G$ be a graph of sufficiently large order $n$. Let $X$ be a non-negative eigenvector corresponding to $\rho(G)$, and $x_{u^{*}}=\max_{u\in V(G)}x_{u}$. If $G$ has the property $P_{t,r}$ and $\rho(G)\geq\sqrt{\lfloor t\rfloor(n-\lfloor t\rfloor)}$, then there exists a vertex subset $V$ of $G$ with $|V|=\lfloor t\rfloor$ such that $x_{v}\geq(1-\frac{1}{50s^{2}})x_{u^{*}}$ and $d_{G}(v)\geq(1-\frac{1}{25s^{2}})n$ for every $v\in V$.
@@
Given that $\rho(G)\geq\sqrt{\lfloor t\rfloor(n-\lfloor t\rfloor)}$ and an arbitrary vertex $u\in V(G)$, we have
@@
Note that $N(u)=A_{1}\cup B_{1}$. Then $N(v)\setminus\{u\}=N_{A_{1}\cup A_{2}}(v)\cup N_{B_{1}\cup B_{2}}(v)$ for each $v\in N(u)$. By the definition of $V_{1}$, it follows that $x_{w}<(5s)^{-4}x_{u^{*}}$ for every $w\in B_{1}\cup B_{2}$. Therefore, we can establish an upper bound for the vertices in $A_{1}$ as follows:
@@
Proof of Theorem 2.1: We start by proving lower bounds on $x_{u}$ and $|N(u)|$ for each vertex $u\in V$. Suppose, for the sake of contradiction, that there exists $u_{0}\in V$ such that $x_{u_{0}}<(1-\frac{1}{50s^{2}})x_{u^{*}}.$
@@
To prove that $|V|=\lfloor t\rfloor$, we begin by assuming that $|V|\geq\lfloor t\rfloor+1$. Based on the previous inequality, every vertex $u\in V$ has at most $\frac{n}{(5s)^{2}}$ non-neighbors. Hence, any $\lfloor t\rfloor+1$ vertices in $V$ have at least $n-(\lfloor t\rfloor+1)\frac{n}{(5s)^{2}}$ common neighbors, where $n-(\lfloor t\rfloor+1)\frac{n}{(5s)^{2}}\geq s$ as $s>t+1$. Then $G$ contains a subgraph isomorphic to $K_{\lfloor t\rfloor+1,s}$. By a similar argument to that in the proof of Claim 2.4, there is a contradiction since $K_{\lfloor t\rfloor+1,s}$ does not satisfy the property $P_{t,r}$. Therefore, $|V|\leq\lfloor t\rfloor.$
@@
Combining both parts, we establish the desired results: $x_{u}\geq\big{(}1-\frac{1}{50s^{2}}\big{)}x_{u^{*}}$ and $|N(u)|\geq\big{(}1-\frac{1}{25s^{2}}\big{)}n$ for each vertex $u\in V$. Moreover, $|V|=\lfloor t\rfloor$. Thus, we complete the proof of Theorem 2.1. ∎
@@
Choose a vertex $u\in U\cup W$. By Claim 3.1, let $N_{U}(u)=\{u_{1},u_{2},\dots,u_{d}\}$ and denote by $H$ the subgraph in $G$ induced by the set $V\cup N_{U}(u)\cup\{u\}$. Then $e(H)\geq(\lfloor t\rfloor+1)d$. Since $H$ has the property $P_{t,r}$, we have $e(H)\leq t|H|+r=t(\lfloor t\rfloor+1+d)+r$. Combining the two inequalities, we have $(\lfloor t\rfloor+1-t)d\leq t(\lfloor t\rfloor+1)+r$, and so
$d_{U}(u)=d\leq\frac{t(\lfloor t\rfloor+1)+r}{\lfloor t\rfloor+1-t}\leq s-1.$ (9)
for each vertex $u\in U\cup W.$
Note that $W$ is not complete to $V$, meaning each vertex in $W$ is not adjacent to some vertex in $V$. By Theorem 2.1, every vertex in $V$ is not adjacent to at most $\frac{n}{25s^{2}}$ vertices of $G$. Recall that $s>t+1$. Therefore,
$|W|\leq\frac{\lfloor t\rfloor n}{25s^{2}}\leq\frac{n}{25s}.$ (10)
Note that $d(u)\leq d_{U}(u)+|V|+d_{W}(u)\leq s-1+\lfloor t\rfloor+d_{W}(u)$. Moreover, since $G$ has the property $P_{t,r}$, either $e(W)\leq t|W|+r$ or $|W|<t$. Consequently, we get that
@@
Suppose that $W\neq\emptyset$. Let $G^{*}$ denote the graph obtained from $G$ by removing all edges within $W$ and from $W$ to $U$, then adding all possible edges from $W$ to $V$. Additionally, let $w_{1}$ be a vertex of minimal degree in $G[W]$, and let $E_{1}$ be the set of edges incident to $w_{1}$ in $G[W]$. Following the same approach, let $w_{2}$ be a vertex of minimal degree in $G[W\setminus\{w_{1}\}]$, and let $E_{2}$ be the set of edges incident to $w_{2}$ in $G[W\setminus\{w_{1}\}]$. Repeating the process, we define $W_{1}=W$ and $W_{i}=W\setminus\{w_{1},\dots,w_{i-1}\}$ for $i\geq 2$, where $w_{i}$ is a vertex of minimum degree in $G[W_{i}]$ and $E_{i}$ is the set of edges that are incident to $w_{i}$ in $G[W_{i}]$. By continuing this process iteratively, we have
@@
In view of the definition of $W$, we have $|V\setminus N_{V}(w_{i})|\geq 1$ for each $w_{i}\in W$. By Theorem 2.1, $x_{v}\geq\big{(}1-\frac{1}{50s^{2}}\big{)}x_{u^{*}}$ for each $v\in V$. Moreover, by inequality (9), we know that $d_{U}(u)\leq s-1$ for each $u\in U$, and based on the previous discussion, we have $d_{W_{i}}(w_{i})<2s$ for each $w_{i}\in W_{i}$. Combining (12) and (3), we can see that
@@
Based on the maximality of $\rho(G)$, the new graph $G^{*}$ cannot satisfy the property $P_{t,r}$. Specifically, there exists an induced subgraph $H$ of $G^{*}$ such that $|H|\geq t$ and $e(H)>t|H|+r$. Since $G^{*}[U\cup V]=G[U\cup V]$, the violation of the property $P_{t,r}$ must involve vertices in $W$, i.e., $V(H)\cap W\neq\emptyset$.
Let $H^{\prime}=G^{*}[V(H)\setminus W]$. According to the definition of $G^{*}$, we know that $N_{G^{*}}(w)=V$ for each $w\in W$. Hence, $e(H)=e(H^{\prime})+|V(H)\cap W|\cdot|V(H)\cap V|.$ Since $|V(H)\cap V|\leq\lfloor t\rfloor\leq t$ and $e(H)>t|H|+r$, it follows that
$e(H^{\prime})>t|V(H)|+r-t|V(H)\cap W|=t|V(H^{\prime})|+r.$
If $|V(H^{\prime})|\geq t$, the above inequality leads to a contradiction because $G^{*}[V(H^{\prime})]=G[V(H^{\prime})]$ and $G$ is assumed to satisfy the property $P_{t,r}$.
If $|V(H^{\prime})|<t$, then by Claim 3.1 and the inequality in (10), we have $|U\setminus V(H^{\prime})|\geq\frac{n}{2}-t\geq|V(H)\cap W|$. Since $U\setminus V(H)=U\setminus V(H^{\prime})$, we can choose a subset $U^{\prime}$ of $U\setminus V(H)$ with $|U^{\prime}|=|V(H)\cap W|$. Denote $H^{\prime\prime}=G^{*}[U^{\prime}\cup(V(H)\setminus W)]$. Then $|V(H^{\prime\prime})|=|V(H)|$, and $N_{G^{*}}(w)=V\subseteq N_{G^{*}}(u)$ for every $w\in W$ and every $u\in U^{\prime}$. Thus, $e(H^{\prime\prime})\geq e(H)>t|V(H)|+r=t|V(H^{\prime\prime})|+r$. This also leads to a contradiction.
@@
By Claim 3.2, we have $V(G)=V\cup U$ and $e(V,U)=|V||U|$, where $|V|=\lfloor t\rfloor$. Based on (9), every $u\in U$ has at most $s-1$ neighbors in $U$. Let $u^{\star}\in U$ such that $x_{u^{\star}}=\max_{u\in U}x_{u}$. Then $\rho x_{u^{\star}}\leq(s-1)x_{u^{\star}}+\lfloor t\rfloor x_{u^{*}}.$ It follows that $x_{u^{\star}}\leq\frac{\lfloor t\rfloor}{\rho-s+1}x_{u^{*}}.$ Also, since $u^{*}\in V$, we have $\rho x_{u^{*}}\leq(\lfloor t\rfloor-1)x_{u^{*}}+(n-\lfloor t\rfloor)x_{u^{\star}}.$ Combining these two inequalities, we get that
@@
4 Characterization of extremal graphs
In this section, we will present the proof Theorem 1.3. Before delving into our proof, we introduce some concepts and several lemmas. An $\ell$-walk consists of $\ell+1$ vertices where consecutive vertices are adjacent. Let $w_{G}^{\ell}(u)$ be the number of $\ell$-walks starting at vertex $u$ in $G$. The total number of $\ell$-walks in $G$ is defined as $W_{\ell}(G)=\sum_{v\in V(G)}w_{G}^{\ell}(v)$. Obviously, $W_{1}(G)=2e(G)$ and $W_{2}(G)=\sum_{v\in V(G)}(d_{G}(v))^{2}.$
Let $G_{1},G_{2}$ be two graphs, potentially of different orders. We say that $G_{1}$ is walk-equivalent to $G_{2}$ and denote this as $G_{1}\equiv G_{2}$, if $W_{\ell}(G_{1})=W_{\ell}(G_{2})$ for all $\ell\geq 1$. If there exists an $\ell$ such that $W_{\ell}(G_{1})>W_{\ell}(G_{2})$ and $W_{i}(G_{1})=W_{i}(G_{2})$ for $i\leq\ell-1$, we say that $G_{1}$ is strictly walk-preferable to $G_{2}$ and denote this as $G_{1}\succ G_{2}$.